# XtalOpt Version 14: Variable-Composition Crystal Structure Search for Functional Materials Through Pareto Optimization

### v2

@@ 13.6 Backwards compatibility
Version 14 of XtalOpt does not read in the input file (xtalopt.in) of previous versions due to various changes in the required input flags, and does not resume a run that was performed with an older version of the code. However, old versions of the code can be readily obtained from the XtalOpt website [78] and the CPC program library.
@@ Appendix A New implementation in the GUI
Figure 4: The graphical user interface (GUI) of XtalOpt version 14 has been updated to include the new features: (a) the “Structure Limits” tab where the chemical system, search type, reference energies, maximum number of atoms, various unit cell limits and cell initialization schemes for the run can be set; (b) the “Search Settings” tab entries can be used to specify the search duration and run concurrency, action taken when a local optimization fails, initial seed structures, genetic operation specifications, and thresholds for symmetry and similarity detections; (c) the “Multiobjective Search” tab parameters allow for adjusting the global optimization type and its relevant settings, adding objectives for a MO run, and specifying the code’s actions when an external programs fails to calculate a filtration objective, and (d) the “Progress” tab maintains a live overview of the produced structures with a summary of their characteristics, such as the chemical formula, enthalpy, distance above the convex hull, Pareto front, symmetry, and ancestry information.
@@ Appendix B Summary of the new, modified, and obsolete CLI flags
New flag Comments (and *default values, if any) Runtime adjustable
@@
Old flag Renamed/changed flag Comments (and *default values, if any) Runtime adjustable

### v3

@@ 13.7 Backwards compatibility
Version 14 of XtalOpt does not read the input file (xtalopt.in) of previous versions due to various changes in the required input flags, and does not resume a run that was performed with an older version of the code. However, old versions of the code can be readily obtained from the XtalOpt website [79] and the CPC program library.
@@ Appendix A New implementation in the GUI
Figure 4: The graphical user interface (GUI) of XtalOpt version 14 has been updated to include the new features: (a) the “Structure Limits” tab where the chemical system, search type, reference energies, maximum number of atoms, various unit cell limits and cell initialization schemes for the run can be set; (b) the “Search Settings” tab entries can be used to specify the search duration and run concurrency, action taken when a local optimization fails, initial seed structures, genetic operation specifications, and thresholds for symmetry and similarity detections; (c) the “Multiobjective Search” tab parameters allow for adjusting the global optimization type and its relevant settings, adding objectives for a MO run, and specifying the code’s actions when an external program fails to calculate a filtration objective, and (d) the “Progress” tab maintains a live overview of the produced structures with a summary of their characteristics, such as the chemical formula, enthalpy, distance above the convex hull, Pareto front, symmetry, and ancestry information.
@@ Appendix B Summary of the new, modified, and obsolete CLI flags
New flag Comments (default values indicated by “*”, if any) Runtime adjustable
@@
Old flag Renamed/changed flag Comments (default values indicated by “*”, if any) Runtime adjustable
@@
Appendix C Example input files for an XtalOpt run
XtalOpt input file for a VC search for the Ti-O system, using the MACE potential.
The content of the xtalopt.in file
softExit = true
chemicalFormulas = Ti7 O1, Ti1 O7
referenceEnergies = Ti3 -23.35901499 , O16 -78.97000885
vcSearch = true
maxAtoms = 40
numInitial = 50
parentsPoolSize = 20
limitRunningJobs = true
runningJobLimit = 2
continuousStructures = 15
maxNumStructures = 500
usingRandSpg = true
minVolumeScale = 3.0
maxVolumeScale = 6.0
weightCrossover = 35
weightStripple = 25
weightPermustrain = 25
weightPermutomic = 15
weightPermucomp = 5
randomSuperCell = 10
radiiScalingFactor = 0.5
aMin = 2.0
bMin = 2.0
cMin = 2.0
aMax = 30.0
bMax = 30.0
cMax = 30.0
alphaMin = 60.0
betaMin = 60.0
gammaMin = 60.0
alphaMax = 120.0
betaMax = 120.0
gammaMax = 120.0
templatesDirectory = ./templates
queueInterface = local
localWorkingDirectory = ./local
numOptimizationSteps = 1
optimizer = vasp
exeLocation = /Users/sam/TiO/vasp_uip.sh
incarTemplates = incar1
kpointsTemplates = kpoints1
potcarFile system = ./templates/potcar1
jobFailLimit = 1
jobFailAction = kill
rdfTolerance = 0.95
spglibTolerance = 0.01
Bash executable script to run the UIP interface script; to be used as a “VASP”-like optimizer for XtalOpt run.
The content of vasp_uip.sh script
#!/bin/bash
# Path to the UIP interface script
exe=/usr/local/bin/vasp_uip.py
# Load python environment
source /Users/sam/penv/bin/activate
# Relaxation task
python \$exe -n 75 -o
With the above files, and with a templates/ folder that contains (empty) files incar1, kpoints1, and potcar1, one can run XtalOpt as:
/path/to/xtalopt --cli
After the run is finished, the pycxl script can be used to plot the convex hull of the run, e.g.:
python /path/to/pycxl.py -i local/hull.txt -y -4 1 -c 1
This produces the out_distances.pdf file, for which an example output is illustrated in Figure 5.
Figure 5: An example of the convex hull produced by the pycxl code, from the output of the XtalOpt VC run for the Ti-O system.